OpenStudy (anonymous):

Suppose the linear cost function C(x) = 6x gives the cost for buying x items. If the items are sold in packages of 10, and no one can buy more than 5 packages, then the RANGE of the function C is

OpenStudy (kropot72):

The range of x is:\[10\le x \le 50\] The range of C is:\[60\le C \le 300\]